OpenStudy (anonymous):

Three times the sum of 4 and a number y squared is the same as the difference of 14and the number y

OpenStudy (anonymous):

Simplify

OpenStudy (anonymous):

can you answer it

OpenStudy (anonymous):

Yep - give me a sec :)

OpenStudy (anonymous):

\[[3(4+y)]^2 = 14-y\] \[(12+3y)^2 = 14-y\] (12+3y)(12+3y) = 14-y 144 + 36y + 36y + 9y^2 = 14 - y

OpenStudy (anonymous):

You then group the like terms, and solve the quadratic (I'll give you a second to try, while I'm writing out the rest)

OpenStudy (anonymous):

130 + 73y + 9y^2 = 0

OpenStudy (anonymous):

Have you been taught to use the quadratic equation to solve these kinds of questions?

OpenStudy (anonymous):

i read it as 3(4+y^2)=14-y 12+3y^2=14-y 3y^2+y+12-14=0 3y^2+y-2=0

OpenStudy (anonymous):

Ah OK! Then you are correct in your working out :) It depends on the punctuation in the question :)

OpenStudy (anonymous):

(The answers for y are y = 2/3 , or y = -1, by the way :) )